### v1

# A Survey on Collaborating Small and Large Language Models for Performance, Cost-effectiveness, Cloud-edge Privacy, and Trustworthiness

Abstract
Large language models (LLMs) have advanced many domains and applications but face high fine-tuning costs, inference latency, limited edge deployability, and reliability concerns. Small language models (SLMs), compact, efficient, and adaptable, offer complementary remedies. Recent work explores collaborative frameworks that fuse SLMs’ specialization and efficiency with LLMs’ generalization and reasoning to meet diverse objectives across tasks and deployment scenarios. Motivated by these developments, this paper presents a systematic survey of SLM–LLM collaboration organized by collaboration objectives. We propose a taxonomy with four goals: performance enhancement, cost-effectiveness, cloud–edge privacy, and trustworthiness. Within this framework, we review representative methods, summarize design paradigms, and outline open challenges and future directions toward efficient, secure, and scalable SLM–LLM collaboration.
A Survey on Collaborating Small and Large Language Models for Performance, Cost-effectiveness, Cloud-edge Privacy, and Trustworthiness
Fali Wang, Jihai Chenthanks:  Work done as an intern at Penn State., Shuhua Yang, Ali Al-Lawati, Linli Tang Pennsylvania State University University Park, USA
Hui Liu Michigan State University East Lansing, USA          Suhang Wangthanks:  Corresponding author. Email: szw494@psu.edu Pennsylvania State University University Park, USA
1 Introduction
Large language models (LLMs) have profoundly transformed multiple domains, including AI for science Luo et al. (2022); Al-Lawati et al. (2025); Wang et al. (2024a), programming Shi et al. (2024), and human-centered interaction Zhang et al. (2024c), primarily owing to their massive parameter scales. However, such scale also introduces several challenges: (1) fine-tuning is computationally intensive, limiting efficient model adaptation Thawakar et al. (2025); Liu et al. (2024b); (2) large model size leads to inference latency, constraining real-time applications Leviathan et al. (2023); Kwon et al. (2023); (3) typical edge devices, such as mobile phones, personal computers, and small servers, lack the capacity to host these models, while cloud-based inference raises privacy and cost concerns Carlini et al. (2021); Xu et al. (2024b); and (4) LLMs exhibit inherent reliability risks, including hallucinations and jailbreak vulnerabilities Yao et al. (2024); Farquhar et al. (2024). These challenges underscore the increasing need for customizable, cost-efficient, edge-deployable, and trustworthy AI solutions. Small language models (SLMs), characterized by their compact architecture, low computational cost, and adaptability, have emerged as promising counterparts to mitigate these issues, although their general reasoning and knowledge coverage remain more limited than those of LLMs. Consequently, harnessing the complementary strengths of SLMs and LLMs provides a viable pathway toward developing efficient, scalable, and reliable AI systems.
Researchers have proposed diverse approaches for SLM–LLM collaboration, leveraging SLMs’ strengths in customization, efficiency, and local deployment alongside LLMs’ generalization and reasoning capabilities Xu et al. (2024a); Chen et al. (2024); Wang et al. (2025b). Despite notable progress, a systematic survey focused on the objectives of SLM–LLM collaboration remains absent. Existing studies largely pursue four objectives. (1) Performance: integrating domain-specific SLMs with general LLMs to improve performance across specialized and general tasks. (2) Cost-effectiveness: using SLMs for lightweight processing and invoking LLMs selectively to reduce computational and API costs. (3) Cloud–edge privacy: employing on-device SLMs for private data handling while cloud LLMs offer broader reasoning to balance efficiency and privacy. (4) Trustworthiness: adopting SLMs as safety policy encoders that guide LLMs to produce safer and more reliable outputs. This paper presents the first comprehensive survey around the objectives of SLM–LLM collaboration: performance, cost-effectiveness, cloud–edge privacy, and trustworthiness. We propose a taxonomy, summarize representative works (Tab.1 in Appendix), and outline future directions.
Differences from existing surveys
Several surveys have examined SLMs in the LLM era. Wang et al. (2024b, 2025a) provided broad overviews of SLM design, applications, and reliability but mentioned collaboration only briefly. Lu et al. (2024); Van Nguyen et al. (2024); Xu et al. (2024b) analyzed SLM advantages, architectures, and deployment, yet offered limited discussion of collaboration. Existing collaboration surveys, such as Chen and Varoquaux (2024); Niu et al. (2025); Li et al. (2025b), focus on one-way or cloud–edge settings without addressing broader objectives. To fill this gap, we systematically review SLM–LLM collaboration by collaboration objectives, offering key insights to guide future studies.
2 Collaboration for Performance
The optimal LLM varies across tasks and queries Jiang et al. (2023), making single-model solutions suboptimal. Leveraging complementary models is therefore effective. Owing to their customizability, SLMs have produced many domain-specific products, such as SciGLM Zhang et al. (2024a), ChemLLM Zhang et al. (2024b), and Biomistral Labrak et al. (2024). Hence, SLM–LLM collaboration for performance has been a prevalent strategy. We divide existing works mainly into two paradigms: (1) Guidance–Generation, where one model steers another’s generation, and (2) Division–Fusion, where the SLMs and LLMs leverage their complementary strengths on distinct tasks.
2.1 Guidance–Generation Collaboration
Guidance–Generation collaboration involves one assistant model providing guidance based on its strengths to support a backbone model in generation. When one model performs well overall and another excels in a specific aspect, the former serves as the generator while the latter offers complementary guidance. This paradigm generally appears in two forms: (1) LLM-guided SLM generation and (2) SLM-guided LLM generation.
Figure 1: Guidance-Generation Collaboration.
(1) LLM-guided SLM generation.
As shown in Fig. 1(a), the LLM uses its semantic understanding to clarify tasks and provide fine-grained guidance for task-specific SLMs. For example, SynCID Liang et al. (2024) employs LLM-generated task descriptions to guide SLM reasoning.
(2) SLM-guided LLM generation.
As shown in Fig. 1(b), the SLM offers domain expertise and cues, while the LLM integrates this information for more accurate and reliable outputs. For instance, SuperICL and its variant Xu et al. (2024a); Mojarradi et al. (2024) inject SLM predictions and confidence scores into the LLM’s context, and LM-Guided CoT Lee et al. (2024) uses SLM-generated reasoning chains to guide LLM inference.
2.2 Division–Fusion Collaboration
Division–fusion collaboration paradigm is preferred when multiple models exhibit heterogeneous capabilities and no single backbone model is suitable. According to different task needs, this paradigm mainly includes (1) Parallel Ensemble for collective intelligence and (2) Sequential Cooperation for multi-stage workflows.
(1) Parallel Ensemble: Collective Intelligence
As shown in Fig. 2, multiple SLMs and LLMs work in parallel, and their outputs are integrated for higher accuracy. Two common strategies are majority voting, as in ELM Gondara et al. (2025), which aggregates outputs by consensus, and cross-verification, as in CaLM Hsu et al. (2024), which iteratively refines results until consensus is reached.
Figure 2: Parallel Ensemble: Collective Intelligence
(2) Sequential Cooperation: Division of Labor
For multi-stage tasks, subtasks are assigned to suitable models and connected in a pipeline, as shown in Fig. 3. When tasks are explicitly staged, SLMs and LLMs take different roles: SLMs handle precise components (e.g., schema matching in ZeroNL2SQL Fan et al. (2024) and KDSL Chen et al. (2025a)), while LLMs manage complex reasoning (e.g., GCIE Bao and Wang (2024)). When stages are implicit, the LLM acts as a planner and SLMs as executors, as in HuggingGPT Shen et al. (2023) and TrajAgent Du et al. (2024), which assign subtasks to expert SLMs and integrate them.
Figure 3: Sequential Cooperation: Division of Labor
2.3 Discussion
Current SLM–LLM collaborations for performance highlight the potential of combining specialized and general models but remain limited by fragmented ecosystems and incomplete evaluation standards. Future research should establish open, cross-domain platforms enabling efficient model discovery and integration, alongside collaborative benchmarks that assess structured cooperation, throughput, and cost–performance trade-offs to better capture real-world system efficiency.
3 Collaboration for Cost-Effectiveness Trade-off
[…]
3.1 Collaboration During Pre-Training Stage
Pre-training is the most computationally intensive phase of the LLM lifecycle, motivating cost-effective strategies. SLM–LLM collaboration improves data efficiency and optimizes training via three approaches: (1) data curation, where LLMs refine SLM data (Fig.4(a)) or SLMs assess LLM data quality (Fig.4(b)); (2) co-curriculum learning (Fig.5(a)), which progressively expands model capacity; and (3) pre-training distillation (Fig.5(b)), transferring knowledge from teacher to student models. (1) Data Curation
(a) SLM selects/Reshape Data and feeds into LLM
(b) Perplexity-based data pruning with SLM
Figure 4: Data Curation.
High-quality data is essential for effective pre-training, yet real-world corpora are often noisy, inflating computational cost and degrading performance. SLM–LLM collaboration addresses this by filtering and refining datasets to improve data efficiency. For example, SALT Rawat et al. (2024) uses compact models to provide soft labels and select informative samples, accelerating LLM pre-training while reducing FLOPs and energy. Conversely, Ankner et al. (2024) employ perplexity-based pruning with SLM scorers to remove low-yield data, reducing optimization steps and enhancing generalization.
(2) Co-Curriculum
(a) Co-Curriculum
(b) Pre-training Distillation
Figure 5: (a) Co-curriculum learning and (b) Knowledge distillation.
LiGO Wang et al. (2023) maps small-model parameters to larger ones via a linear growth operator, cutting pre-training compute by up to 50%. Progressive training Yano et al. (2025) incrementally scales width and depth, saving $\sim$25% compute. HyperCloning Samragh et al. (2024) clones linear layers for stable warm starts and substantial GPU-hour reductions.
(3) Pre-Training Distillation Pre-training distillation (PD) extends classic task-agnostic methods such as DistilBERT, TinyBERT, and MiniLM Sanh et al. (2019); Jiao et al. (2020); Wang et al. (2020), now scaled to LLMs. Recent studies show that larger students can sometimes outperform teachers Peng et al. (2025) and disentangle PD effects from data confounds Goyal et al. (2025). MINIPLM Gu et al. (2025) refines data with teacher inference before student training, enhancing efficiency and diversity.
3.2 Collaboration During Tuning Stage
LLM tuning is highly resource-intensive, demanding substantial GPU memory and FLOPs, while SLM tuning distilled from LLMs can suffer from inefficient or low-quality supervision, leading to wasted LLM computation. SLM–LLM collaboration alleviates these by relocating supervision to low-cost venues and compressing guidance: (1) Selective Knowledge Distillation (SLM $\Leftarrow$ LLM, Fig. 6) distills selective LLM signals to reduce LLM calls; and (2) Proxy Transfer (SLM $\Rightarrow$ LLM, Fig. 7(a), Fig. 7(b)) generates lightweight, reusable updates for efficient LLM adaptation.
(1) Selective Knowledge Distillation.
Figure 6: Selective Knowledge Distillation.
Adaptive selection reduces data and teacher calls (LLKD Li et al. (2025a), efficiency-oriented KD Jazbec et al. (2024)); dense supervision uses concise yet rich signals such as contrastive label refinement (SynCID Liang et al. (2024)), compact CoT (HAWKEYE She et al. (2025)), and causal explanations Muhebwa and Osman (2025); targeted matching focuses on task-relevant features (C2KD Chen et al. (2025c)). Behavioral KD distills tool-use trajectories (Agent Distillation Kang et al. (2025)). Applications, e.g., LSC4Rec Lv et al. (2025), SmallPlan Pham et al. (2025), and Think Bergner et al. (2024), show that these strategies jointly cut FLOPs, tokens, and LLM calls while retaining near-LLM performance under tight compute and latency constraints.
(2) Proxy Transfer
(a) SLM-trained Compatible Parameters
(b) SLM-output Merging
Figure 7: Proxy Transfer.
Models can train parameter updates on a small proxy model and lift them for use in the LLM. LoRAM Zhang et al. (2025a) trains low-rank adapters on an SLM to locate effective subspaces, later reused in LLMs. LiteMOE Zhuang et al. (2024) extracts proxy submodels for on-device tuning, enabling local specialization and feedback propagation. GateKeeper Rabanser et al. (2025) cascades an SLM and LLM to handle easy cases locally and defer harder ones, cutting inference cost, while logit-fusion Fan et al. (2025c) integrates fine-tuned SLM outputs to refine LLM distributions with minimal updates.
3.3 Collaboration During Inference Stage
Inference cost arises from three factors: (i) executing large models for every query (high FLOPs, latency, and API cost), (ii) decoding inefficiency (token-by-token verification by a single heavy model), and (iii) misallocated compute (overspending on easy queries, underspending on hard ones). SLM–LLM collaboration mitigates these through: (1) Cascade routing, handling easy queries with an SLM and escalating difficult ones (Fig. 8); (2) Speculative decoding, an SLM drafts tokens that a stronger verifier approves in bulk (Fig. 9); and (3) Compute-optimal test-time scaling, optimizing model selection and sampling budgets across SLM/LLM under fixed budget (Fig. 10).
(1) Cascade Routing
Figure 8: Cascade Routing
Query-level routers route per input using lightweight difficulty predictors Chen et al. (2024); Ong et al. (2025); Zhang et al. (2024d); Yue et al. (2024); Hao et al. (2024), e.g., SlimPM flags missing LLM knowledge for retrieval Tan et al. (2024). Token-level routers escalate only complex spans to LLMs, improving cost–quality trade-offs Zheng et al. (2024); Fu et al. (2025a). Resource-aware routers optimize selection via contextual bandits Wang et al. (2025c) or MoE-style edge–cloud load balancing Jin et al. (2025).
(2) Speculative Decoding
Figure 9: Speculative Decoding
Speculative decoding advances focus on (i) decoupled draft–verify architectures, (ii) parallel decoding, and (iii) efficient deployment topologies. Cross-attention drafters (Beagle Zhong et al. (2025)) simplify verifier fusion; LookAhead Reasoning Fu et al. (2025b) and Adaptive Parallel Decoding Israel et al. (2025) improve parallelism; edge–cloud systems (LLMCAD Xu et al. (2023), PICE Zhan et al. (2025), FS-GEN Zhang et al. (2024e)) reduce verification cost via local drafting and adaptive policies.
(3) Compute-optimal Test-time Scaling
Figure 10: Compute-optimal TTS
Test-Time Scaling (TTS) increases inference compute to improve performance, while compute-optimal TTS maximizes it under fixed budgets by selecting suitable model sizes and families, representing a form of SLM–LLM collaboration. BestRoute Ding et al. (2025) optimizes model routing and TTS sampling budgets to reduce cost, and AgentTTS Wang et al. (2025b) extends this to multi-stage tasks through dynamic model and budget allocation.
3.4 Discussion
Existing cost-effectiveness collaborations show clear savings across compute, latency, communication, storage and API spend by coordinating the ML lifecycle, but remain constrained by fragmented pipelines, inconsistent cost reporting and brittle rules for when to hand off to the LLM Behera et al. (2025). Future work should standardize end-to-end cost-effectiveness metrics and use adaptive routing–paired with compression, bidirectional distillation etc, to deliver reliable savings across SLM-LLM collaboration paradigms.
4 Privacy-preserving Cloud–Edge Collaboration
Deploying LLMs on edge devices is limited by storage and computation constraints. A practical alternative is a cloud–edge architecture where lightweight SLMs handle local, context-aware tasks, and cloud-based LLMs provide broader generalization. However, transmitting sensitive data from trusted edge SLMs to semi-trusted cloud LLMs raises privacy concerns, as personally identifiable or proprietary information may be logged and leaked for training Duan et al. (2023), posing long-term security and compliance risks.
To balance privacy and performance, privacy-preserving collaboration has emerged at both the inference and fine-tuning stages. At the inference stage, SLMs act as Sensitive Information Gatekeepers, filtering or anonymizing data before sending it to LLMs, or as All-Information Guardians, retaining private context locally and integrating it after generation. At the fine-tuning stage, collaboration occurs through SLMs as Learners (learning from LLMs via distillation), LLMs as Learners (adapting via proxy tuning using SLM signals), or Collaborative Learners (jointly improving through federated knowledge transfer).
4.1 Inference Stage
During inference, edge user requests often contain sensitive information. Edge SLMs can process them locally but with lower quality than cloud LLMs, while direct cloud LLM use risks data exposure. Thus, privacy-preserving edge–cloud collaboration is needed, based on protection levels, categorized into (1) Sensitive Information Gatekeepers and (2) All-Information Guardians.
(a) Sensitive information sanitization
(b) Selective disclosure
Figure 11: SLMs as Sensitive-Information Gatekeepers.
(1) SLMs as Sensitive Information Gatekeepers To protect sensitive information in edge–cloud collaboration, the edge SLM serves as a privacy-preserving filter, for (i) sanitizing sensitive content and (ii) selectively disclosing necessary information before transmitting requests to the cloud.
(i) Sanitizing sensitive information: Fig.11(a) shows how SLMs remove or rephrase prompts containing personally identifiable information (PII). Chong et al. (2024) use an SLM to filter PII before cloud transmission, while Papadopoulou et al. (2022) employ SLM probabilities for detection. Uzor et al. (2025) rephrase queries to preserve meaning while removing PII, applied to clinical data Wiest et al. (2024). Hartmann et al. (2024) further enhance SLMs by masking sensitive data and using LLM-generated in-context examples.
(ii) Selective disclosure: Fig. 11(b) illustrates selective disclosure, where the edge SLM shares limited, task-relevant information to reduce privacy risks. Zhang et al. (2024f) share sketches or logits with the LLM, while CORE Fan et al. (2025a) ranks and filters relevant context. Zhang et al. (2024d) employ a policy network to balance privacy and quality, and Cheng et al. (2025) enforce a differential privacy budget to mathematically bound potential leakage. In sum, these methods shift from full-text disclosure to abstracted representation sharing, mitigating data exposure in SLM–LLM collaboration.
(2) SLMs as All-Information Guardians
Figure 12: SLMs as All-Information Guardians.
Edge devices often enforce strict privacy rules that prevent PII disclosure. In such settings, SLMs serve as All-Information Guardians, handling sensitive data on-device while drawing on cloud LLM knowledge, ensuring privacy preservation (Fig. 12). Sketch-based CoGenesis Zhang et al. (2024f) illustrates this by using cloud LLMs for planning and local SLMs for personalized adaptation. Similar frameworks integrate cloud knowledge with local data for clinical Zhang et al. (2024g) or retrieval-augmented Cheng et al. (2025); Hartmann et al. (2024) applications. Industrial systems such as Apple Intelligence Gunter et al. (2024) follow the same principle, processing sensitive data locally and sharing only abstractions with the cloud.
4.2 Fine-tuning Stage
During fine-tuning, edge data can enhance both edge and cloud LLMs, though protecting sensitive information remains challenging. Depending on the enhanced target, privacy-preserving edge–cloud frameworks typically adopt three paradigms: (1) SLMs as Learners, gaining knowledge from LLMs through local fine-tuning or distillation; (2) LLMs as Learners, improving via proxy signals from tuned SLMs; and (3) Collaborative Learners, where both models co-evolve through federated or privacy-preserving knowledge exchange.
(1) SLMs as Learners
Figure 13: SLMs as Learners.
Fig. 13 shows edge SLMs learning from cloud LLMs to improve their ability and reduce cloud dependence through subsequent local fine-tuning. SLMs can learn via two modes: (i) raw text, such as reasoning traces or samples, and (ii) feedback logits as supervision signals.
(i) Raw Data: SLMs enhance reasoning by learning from cloud LLM outputs while keeping private data local. The LLM provides interpretable supervision (e.g., responses, reasoning traces) or synthetic samples for SLM fine-tuning. MiniLLM Gu et al. (2024) and LlamaDuo Park et al. (2025) distill reasoning traces, while DRAG Chen et al. (2025b) extends to retrieval-augmented reasoning. Privacy-aware frameworks such as HomeLLaMA Huang et al. (2025b) and Qin et al. (2024) use selective textual supervision to align SLMs with user preferences under strict privacy constraints.
(ii) Logit-based Learning: SLMs learn from teacher LLMs’ numerical outputs, using logits as soft targets to imitate their decision boundaries without accessing raw text. TinyLLM Tian et al. (2025) distills from multiple teachers for robustness, Mix Distillation Li et al. (2025d) combines LLM and SLM supervision, and ADPA Gao et al. (2025) introduces an advantage-based objective for better reasoning alignment, ensuring privacy and efficiency through abstracted signal exchange.
(2) LLMs as Learners
Figure 14: LLMs as Learners.
LLMs can be adapted to sensitive data without direct access by leveraging an edge SLM tuned locally. The SLM forwards only supervision signals, such as (i) logits or (ii) LoRA signals, to the LLM, enabling privacy-preserving alignment, as shown in Fig. 14.
(i) Logit-based Learning: Instead of directly fine-tuning LLMs, Ormazabal et al. (2023) steer them via SLMs trained on private data, exposing only logits for guidance, with later extensions using logit offsets Liu et al. (2024a); Wang et al. (2025d). Follow-up work addresses training–inference inconsistency He et al. (2024); Yao et al. (2025) and improves robustness through re-encoding SLM outputs Fang et al. (2024a), enabling adaptation across SLM–LLM pairs without shared vocabularies while preserving data privacy.
(ii) LoRA-based Learning: Recent work explores using LoRA signals for knowledge transfer. PHLoRA Vasani et al. (2025) avoids dual logit outputs by computing low-rank adaptations with strong performance gains. LoRASuite Li et al. (2025c) and LoRA-X Farhadzadeh et al. (2025) enable cross-version and cross-architecture migration, while Trans-LoRA Wang et al. (2024c) and Cross-LoRA Xia et al. (2025) generalize LoRA transfer across heterogeneous models for scalable, privacy-conscious LLM adaptation.
(3) SLMs and LLMs as Collaborative Learners
Figure 15: Cloud-Edge Collaborative Learning.
Beyond one-way transfer, collaboration can be bidirectional: SLMs and LLMs co-adapt through shared representations, logits, or adapter weights while preserving privacy (Fig. 15). Edge SLMs train on local data and send abstracted updates (e.g., logits, LoRA signals) for cloud aggregation. CrossLM Deng et al. (2025) combines private data and LLM-generated samples for co-training, while LSRP Zhang et al. (2025b) enables mutual retrieval and response refinement. Federated frameworks Fan et al. (2025b); Gao et al. (2024); Peng et al. (2024); QI et al. (2024); Fang et al. (2024b) aggregate SLM updates to refine central LLMs, with FLoRA Wang et al. (2024d) enhancing personalization through improved edge aggregation.
4.3 Discussion
Current cloud–edge SLM–LLM collaborations face challenges in achieving secure peer-to-peer exchange, effective cold-start learning, formal privacy guarantees, and scalable personalization under strict data protection. Future work should develop adaptive frameworks enabling privacy-preserving collaboration and lightweight personalization for policy-compliant intelligence in sensitive domains such as healthcare and customer service.
5 Collaboration for Trustworthiness
LLMs show strong generative ability but face trustworthiness issues, including hallucinations, hazards, bias, privacy leaks, and jailbreak. Updating large models for safety is costly and inflexible. As a remedy, SLMs offer lightweight, adaptable external safety control. Recent collaborations adopt two forms: (1) Safety-guided decoding, where SLMs adjust LLM logits during generation, and (2) Guardian–generator, where SLMs filter inputs or outputs, ensuring safer LLM deployment.
5.1 Safety-Guided Decoding
Unsafe LLM outputs often emerge during decoding, making token-level safety control essential yet difficult to anticipate. Recent methods use safety-tuned SLMs to guide LLM decoding against jailbreaks, privacy leaks, and bias. As shown in Fig. 16, SLM and LLM logits are fused for safe, high-quality generation via (1) direct logit fusion or (2) offset fusion, where SLM logits steer LLM probabilities toward safer tokens.
Figure 16: SLM Guided Safe LLM Decoding.
(1) Direct Logit Fusion The CP-$\Delta$ theory Vyas et al. (2023) provides an information-theoretic foundation for model fusion, demonstrating that combining the predictive distributions of two models, $q_{1}(y|x)$ and $q_{2}(y|x)$, with distinct copyright risks yields a fused distribution $p(y|x)\propto q_{1}(y|x)^{\alpha}q_{2}(y|x)^{1-\alpha}$ that minimizes weighted KL divergence and balances their strengths without retraining, where $\alpha$ controls each model’s contribution. The fused distribution lies closer to both sources than they are to each other, thus achieving a purification effect, retaining desirable properties while reducing copyrighted content. This operation is equivalent to linear weighting in logit space, $z_{p}=\alpha z_{1}+(1-\alpha)z_{2}$, providing the theoretical foundation for SLM–LLM collaborative logit fusion. Subsequent methods build on this principle: Purifying LLMs Li et al. (2024) fuses clean SLM and contaminated LLM logits to mitigate copyright, privacy, and backdoor risks, while MOD Shi et al. (2024) extends fusion to multiple objectives, integrating models optimized for fluency, faithfulness, and safety to enable controllable generation.
(2) Logit Offset Fusion Mitchell et al. (2024) introduced the concept of logit offset fusion by decomposing a fine-tuned model’s output into the base model’s logits and a behavioral delta. Through an up-scaling strategy, it integrates a large pre-trained model’s logits with the behavioral delta of a small fine-tuned model, achieving 73% of full large-model fine-tuning helpfulness without retraining. This demonstrates that safety-aligned small models’ behavioral deltas can enhance LLM safety when fused at the logit level. Building on this idea, Offset Unlearning Huang et al. (2025a) trains SLMs to generate targeted logit offsets that suppress sensitive or fabricated knowledge for privacy-preserving unlearning. Conversely, Weak-to-Strong Jailbreak Zhao et al. (2025) reveals that unsafe SLMs can steer safe LLMs toward harmful outputs, implying that the same mechanism can defensively employ safe SLMs to counter jailbreak attacks through safety-guided logit fusion.
5.2 SLM as Guardian, LLM as Generator
Guardrail-based approaches enhance LLM safety by placing SLMs before and after generation. In the SLM-as-Guardian and LLM-as-Generator paradigm, as shown in Fig. 17, SLMs, encoding the safety policy, handle input filtering and output auditing, offering configurable, interpretable, and reusable safety control across evolving applications.
Figure 17: SLM Guardian, LLM Generator
Building on the SLM-as-Guardian paradigm, subsequent works further advance safety coverage and efficiency. The Llama Guard series Inan et al. (2023); Team (2024); Dubey et al. (2024); Fedorov et al. (2024); Chi et al. (2024); AI (2025a) evolves from basic LLaMA-based classifiers to multilingual, cross-modal, and domain-adapted guardrails with fine-grained risk taxonomies. Lightweight filters such as Prompt Guard AI (2024, 2025b), ShieldGemma Zeng et al. (2024, 2025), and WildGuard Han et al. (2024) improve efficiency and robustness. ThinkGuard Wen et al. (2025) enhances interpretability, MiniCheck Tang et al. (2024) supports factual verification, Cascade Guardrails Nagireddy et al. (2024) introduce hierarchical defenses, and LlamaFirewall Chennabasappa et al. (2025) enables multi-agent protection.
5.3 Discussion
Current SLM–LLM collaborations for trustworthiness face challenges in unifying decoding and guarding, extending safety beyond text, enabling edge deployment, and standard evaluation. Future work should build integrated frameworks with multimodal safety, lightweight edge guardians, and standard evaluation benchmarks.
6 Challenges and Future Directions
More future directions are in Appendix A.2.
Open and Interoperable Ecosystem. Existing SLM–LLM collaborations are fragmented and task-specific. Future work should build open, cross-domain ecosystems with standardized interfaces and shared repositories to enable model reuse, composition, and joint benchmarking.
Comprehensive Benchmarking. Current evaluations emphasize open-ended performance but neglect structured and system-level metrics. Future benchmarks should assess structured reasoning, modular cooperation, and efficiency indicators such as throughput, latency, and cost–performance.
Unified Safety Collaboration. Present safety methods decouple decoding and moderation. Future research should unify them, allowing SLMs to act as safety-aware decoders and guardians for end-to-end, trustworthy generation and interaction.
7 Conclusions
In this paper, we present a systematic survey of SLM–LLM collaboration from the perspective of collaboration objectives: performance enhancement, cost-effectiveness, cloud–edge privacy, and trustworthiness. We analyze representative methods showing that most leverage SLMs’ lightweight adaptability. We aim to highlight the current research state and provide insights for future work in this promising area of SLM–LLM collaboration.
Limitations
Although our work introduces a systematic framework for analyzing SLM–LLM collaboration, several limitations remain. First, given the rapid evolution of language model technologies, certain implicit forms of SLM–LLM collaboration may not yet be fully captured, and further insights from the research community are encouraged to complement this study. Second, overlaps exist among several collaboration objectives, for instance, the personalization synergy, while not addressed in this survey, often intersects with cloud–edge collaboration, yet these cross-objective studies remain limited despite their conceptual and practical significance. We expect that addressing these challenges will motivate future comprehensive surveys and deeper investigations into SLM–LLM collaboration.
Appendix A Appendix
A.1 Applications
We summarize the representative works in Table 1, detailing their collaboration mechanisms, employed SLMs and LLMs, evaluation datasets, and goals. From these studies, we can summarize the applications under each category.
Collaboration for performance In Guidance–Generation, SLMs aid LLMs in reasoning and classification (intent detection, NLI, sentiment, multi-hop QA). In Division–Fusion, they form pipelines for domains such as clinical analysis, factual verification, NL2SQL, and trajectory modeling, with SLMs handling subtasks and LLMs integrating results.
Collaboration for cost-effectiveness These show stage-specific application preferences. Inference-stage approaches emphasize real-time reasoning and conversational tasks where latency and compute efficiency are critical. Tuning-stage methods focus on intent classification and reasoning transfer, enabling smaller models to inherit LLM capabilities with minimal cost. Pre-training-stage frameworks target corpus construction and scaling efficiency on large datasets, supporting sustainable and scalable model development.
Collaboration for cloud-edge privacy Inference-stage solutions suit end-user scenarios with strict privacy needs, where edge SLMs prevent PII or sensitive context leakage in domains such as healthcare and customer service. Depending on privacy levels, SLMs can act as Sensitive Information Gatekeepers or All-Information Guardians. Meanwhile, proxy tuning enables domain adaptation of cloud LLMs without direct data exposure, offering broader applicability under edge resource limits and SLMs’ constrained reasoning capacity.
Collaboration for trustworthiness SLM–LLM collaboration for trustworthiness has been applied in safety-critical and compliance-sensitive contexts. Safe decoding methods mitigate hallucination, privacy, and toxicity risks in code generation, instruction following, and dialogue. Guardian frameworks support moderation, assistants, and fact verification by filtering unsafe inputs and ensuring output integrity. Together, these works show that SLMs function as lightweight, adaptive safety modules, embedded for fine-grained control or external for interactive oversight, enabling scalable, reliable AI deployment.
A.2 More Future Directions
Open and Interoperable SLM–LLM Ecosystem Current SLM–LLM collaborations remain fragmented, with isolated models and task-specific systems limiting interoperability and reuse. Future research should develop open, cross-domain ecosystems that unify specialized and general models through standardized interfaces and shared repositories, enabling efficient model discovery, seamless composition, and collaborative benchmarking across diverse applications and domains.
Benchmark Design for SLM–LLM Collaboration Current SLM–LLM evaluations focus mainly on open-ended tasks and performance, overlooking structured tasks and system-level metrics. Future research should develop diverse benchmarks covering structured reasoning, modular cooperation, and system-level metrics, such as throughput, latency, and cost–performance, to better capture collaborative efficiency and real-world applicability.
Synergistic SLMs as Decoders and Guardians. Current safety mechanisms often separate decoding control from post-hoc moderation, limiting effectiveness in SLM–LLM safety collaboration. Future research should develop unified frameworks where SLMs function simultaneously as safety-aware decoders and guardians, enabling multi-stage, end-to-end safety governance that enhances trustworthiness across both generation and interaction phases.
Cost-Aware SLM–LLM Collaboration Current SLM–LLM collaboration paradigms mostly prioritize performance gains without a rigorous, comprehensive cost model, potentially leading to inefficient resource allocation where the operational expense negates the utility benefits. Future research can develop intelligent, adaptive orchestration frameworks that dynamically optimize the cost-performance trade-off. This includes investigating predictive cost-aware routing to direct queries to the most economically efficient model, designing non-myopic collaboration policies that consider long-term inference budgets, and creating multi-objective optimization techniques that balance accuracy, latency, and financial cost for sustainable and scalable SLM–LLM systems.
Theory for Privacy-Preserving Collaboration Current privacy-preserving SLM–LLM collaborations lack formal theoretical foundations, relying mostly on empirical evaluations without quantifiable privacy assurances. Future research should establish rigorous metrics and provable bounds for information leakage and privacy preservation, enabling verifiable guarantees of data protection and fostering trust in large-scale, privacy-sensitive SLM–LLM systems.
Table 1: Representative works on collaborations across performance, cost-effectiveness, privacy, and trustworthiness.
Collab. Mechanism Paper SLM LLM Evaluation Dataset Goal / Contribution
I. Performance Collaboration
LLM guided SLM SynCID (2024) BERT GPT-3.5 Banking77, CLINC150 Refine labels for intent detection
SLM guided LLM SuperICL (2024a) RoBERTa GPT-3.5/4 SST-2, MNLI, etc. Inject SLM predictions and confidence
SLM guided LLM Ensem. SuperICL (2024) BERT/RoBERTa GPT-3.5/LLaMA SST-2, MNLI, etc. Multi-SLM guidance for LLMs
SLM guided LLM LM-Guided CoT (2024) Flan-T5 ChatGPT HotpotQA, 2WikiMultiHopQA Provide rationales to improve reasoning
Parallel ELM (2025) ClinicalBERT GPT-4 BCCR Pathology Reports Aggregate SLM votes; LLM verifies
[…]
Sequential ZeroNL2SQL (2024) T5, BART GPT-3.5/4 Spider, WikiSQL, etc. SLM drafts SQL; LLM finalizes
Sequential KDSL (2025a) Qwen2-VL 2B Qwen2-VL 7B SLM rules; LLM inference
Sequential HuggingGPT (2023) CLIP, BLIP, T5 GPT-3.5 ImageNet, COCO, etc. LLM orchestrates multimodal SLMs
Sequential TrajAgent (2024) TrajFormer GPT-4 TaxiBJ, Porto, etc. LLM plans; SLMs execute
II. Cost-effective Collaboration
Cascade R2R (2025a) DeepSeek-R1-1.5B DeepSeek-R1-32B AIME, GPQA Token-level routing for efficiency
Speculative Beagle (2025) LLaMA-68M LLaMA-7B ShareGPT Cross-attention drafting
TTS-optimal AgentTTS (2025b) LLaMA-3-3B LLaMA-3-70B HotpotQA, CWQ, etc. Budget-aware collaboration
Selective KD SynCID (2024) BERT text-davinci-003 BANKING, CLINC Latent-space alignment distillation
Proxy transfer LiteMoE (2024) Gemma-2B Gemma-7B MMLU, GSM8K, etc. Efficient proxy tuning
Data curation SALT (2024) Transformer-1.5B Transformer-2.8B Pile SLM-guided data selection
Co-curriculum Yano et al. (2025) LLaMA-1B LLaMA-8B FineWeb-Ed Joint curriculum design
Pre-train distill. Peng et al. (2025) GLM-4-1.9B GLM-4-9B MMLU, C-Eval, etc. Progressive knowledge transfer
III. Privacy-aware Cloud–Edge Collaboration
PII gatekeeper Uzor et al. (2025) Phi-3.5, Gemma-2 GPT-4o HealthTap, MeQSum Rewrite prompts to remove PII
All-info guardian CoGenesis (2024f) StableLM-Zephyr GPT-4-turbo Avocado Research Email Cloud knowledge for edge inference
All-info guardian Hartmann et al. (2024) Gemini 1.0 Nano-2 Gemini 1.0 Ultra GSM8K Cloud ICL demos for edge SLM
SLM learner MiniLLM (2024) GPT-2-125M GPT-2-1.5B DollyEval, SelfInst, etc. Distill LLM into SLM
LLM learner Proxy-tuning (2024a) LLaMA-2-7B LLaMA-2-70B GSM, Toxigen, etc. Steer LLM with private SLM
Co-learners FDLoRA (2024) LLaMA-2-7B LLaMA-2-7B BGL, Spirit, etc. Federated LoRA with sharing
IV. Trustworthiness Collaboration
Safe decoding MOD (2024) Safety branch LLaMA-2-13B/70B Logit fusion for safety–utility
Safe decoding W2S Jailbreaking (2025) LLaMA-2-13B LLaMA-2-70B AdvBench, MaliciousInstruct Analyze SLM logit effects
Safe decoding Offset Unlearning (2025a) LLaMA-2-7B LLaMA-2-13B TOFU Logit offset for unlearning
Safe decoding LLM Purifying (2024) CodeLLaMA-7B CodeLLaMA-13B HumanEval, LAMBADA, etc. Logit fusion for clean gen.
SLM guard LLaMA Guard (2023) LLaMA-2-7B LLaMA-70B ToxicChat, OpenAI ModEval I/O safety screening
SLM guard Prompt Guard / 2 (2024; 2025b) DeBERTa-v2-22M/86M LLaMA/GPT-family Red-teaming (multilingual), etc. Lightweight input filtering
SLM guard ShieldGemma / 2 (2024; 2025) Gemma-2B/9B/27B Any ToxicChat, OpenAI Mod, etc. Fine-grained moderation
SLM guard STAND-Guard (2024) Mistral-7B GPT-3.5/4 BeaverTails, SafeRLHF Task-adaptive moderation
SLM guard WildGuard (2024) Mistral-7B-v0.3 Any WildGuardMix Multi-risk detection
SLM guard ThinkGuard (2025) LLaMA-3-8B Any BeaverTails, ToxicChat Reasoned moderation
SLM guard MiniCheck (2024) Flan-T5 / UL2 GPT-4 / Claude LLM-AggreFact Fact verification